OpenStudy (anonymous):

Which value of x satisfies both -9x + 4y = 8 and -3x − y = 4?

OpenStudy (anonymous):

– 1 7 – 7 9 – 7 8 – 9 – 4 3

OpenStudy (er.mohd.amir):

multiply by 4 second eq -12x-4y=16 add wit first -9x+4y-12x-4y=8+16 -21x=24 x=-24/21 =-8/7

OpenStudy (anonymous):

-7/8?

OpenStudy (anonymous):

@MrProphecyXII

OpenStudy (anonymous):

sorry did not notice I was Tagged lol

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

its ok bro